Find the pressure at the bottom of a 10 ft high water filled cylinder. (The density of water is 62.4 lb/ft3.)

6.24 lb/ft2

6.24 lb/ft3

62.4 lb-ft2

624 lb/ft2
Answer #1

Pressure at any point in a fluid exerted by fluid above the point is called hydrostatic pressure.

Hydrostatic Pressure, P=hg

Where, h=height of fluid above the point

=density of fluid

g=acceleration due to gravity

Here, h=10 feet

g 10 m/s2=10*3.2808 ft/s2

=​​​​​ 62.4 lb/ft3

Pressure at the bottom of the cylinder,

P=10 * 62.4 *g lbf/ft2​​​​​​

=62.4lbf/ft2 (1 lb= 1 lb * g)

​​​​​​